OpenStudy (anonymous):

11. what is the equation of the line perpendicular to the line y= 4/9x -2 that passes through the point (4,3)?

OpenStudy (anonymous):

An equation perpendicular will have a slope as the opposite reciprocal of the slope to that line.

OpenStudy (anonymous):

A. y= -4/9x-2 B. y= 9/4x+6 C. y= - 9/4x+12 D. y=-9/4x-12

OpenStudy (anonymous):

so -9/4??

OpenStudy (anonymous):

This means, that since your slope is (4/9), the opposite reciprocal slope will be (-9/4), so the equation y = -9/4x - 2 is a line perpendicular to y = 4/9x - 2

OpenStudy (anonymous):

You got it!

OpenStudy (anonymous):

Ok, now, there's one more step.

OpenStudy (anonymous):

It passes through the coordinate (4, 3). So this means that we can say x = 4 and y = 3. Since we have a slope (m) of the perpendicular line as: m = -9/4, we can set up an equation like this Slope-Intercept Form: y = mx + b Substituting x, y, and m 3 = (-9/4)(4) + b And we can solve for b! (Which is the y-intercept)

OpenStudy (anonymous):

3 = (-9/4)(4) + b 3 = -36/4 + b Do you know how to solve for 'b' here?

OpenStudy (anonymous):

3 = -36/4 + b

OpenStudy (anonymous):

Let me know if you are lost, I can go slower.

OpenStudy (anonymous):

12?

OpenStudy (anonymous):

Lemme see,